OpenStudy (anonymous):

log10^2 +16log10^16/25+12log10^25/24+7log10^81/80+log10^25/49+2log10^7 =?

OpenStudy (anonymous):

the ans=2log5+1

OpenStudy (anonymous):

@TuringTest plzz help

OpenStudy (anonymous):

try moving all factors of logs to exponents, then exponentiate and simplify?

OpenStudy (anonymous):

I will work it this way and see where it takes me. :)

OpenStudy (anonymous):

i did but not getting dude plzz can u show me

OpenStudy (anonymous):

When I work it through, I will.

OpenStudy (callisto):

\[log10^2 +16log10^{\frac{16}{25}}+12log10^{\frac{25}{24}}+7log10^{\frac{81}{80}}+log10^{\frac{25}{49}}+2log10^7 \] Is this the question?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is it log base 10 of 10^apower

OpenStudy (anonymous):

but the ans=2log5 + 1

OpenStudy (callisto):

\[log10 =1\]\[logx^a= alogx\] \[log10^2 +16log10^{\frac{16}{25}}+12log10^{\frac{25}{24}}+7log10^{\frac{81}{80}}+log10^{\frac{25}{49}}+2log10^7 \]\[=2log10 +16(\frac{16}{25})log10+12(\frac{25}{24})log10+7(\frac{81}{80})log10+(\frac{25}{49})log10+2(7)log10 \]\[=2 +16(\frac{16}{25})+12(\frac{25}{24})+7(\frac{81}{80})+(\frac{25}{49})+2(7) \] \[Can \ you \ do \ it ? \]

OpenStudy (anonymous):

then the answer wont be 2log5 +1

OpenStudy (callisto):

Oh... then either the question you posted is incorrect, or the answer you've looked at is incorrect...

OpenStudy (anonymous):

oh........ it is log 2 to the base 10 and ............go o n all other too

OpenStudy (callisto):

._. That's totally different......

OpenStudy (callisto):

\[log_{10}2?\]

OpenStudy (anonymous):

yaa

OpenStudy (callisto):

\[log_{10}2 = log2\]

OpenStudy (anonymous):

I am at 10^x=2+(16/25)^16+(25/24)^12+(81/80)^7+25/49+7^2 have you tried this approach yet?

OpenStudy (anonymous):

how log10^2=log2

OpenStudy (anonymous):

log is assumed base ten unless otherwise specified.

OpenStudy (callisto):

Wait is it \[log10^2 \ or \ log_{10}2?\]

OpenStudy (callisto):

log with a base 10 is common log, usually, we omit the base 10... and write log 2 directly

OpenStudy (anonymous):

or......... is the question

OpenStudy (anonymous):

i am unable to write the equation it is not log 10 square the other one is the question

OpenStudy (anonymous):

could you scan the problem and post it as image? just so we are all clear on the problem

OpenStudy (callisto):

\[log2^2 +16log2^\frac{16}{25}+12log2^\frac{25}{24}+7log2^\frac{81}{80}+log2^\frac{25}{49}+2log2^7 \]Is this the question?

OpenStudy (anonymous):

log with a base 10 is common log, usually, we omit the base 10... and write log 2 directly this is correct

OpenStudy (anonymous):

this is the question.

OpenStudy (callisto):

\[log2^2 +16log2^\frac{16}{25}+12log2^\frac{25}{24}+7log2^\frac{81}{80}+log2^\frac{25}{49}+2log2^7 \]\[=2log2 +16(\frac{16}{25})log2+12(\frac{25}{24})log2+7(\frac{81}{80})log2+(\frac{25}{49})log2+2(7)log2 \]\[=[2 +16(\frac{16}{25})+12(\frac{25}{24})+7(\frac{81}{80})+(\frac{25}{49})+2(7) ]log2 \]Can you simplify it first and tell me what you've got?

OpenStudy (callisto):

Anyway, It won't give the answer you provided lol

OpenStudy (anonymous):

u got the question wrong

OpenStudy (anonymous):

log10^2 the 10 is base and 2 is the exponent eg log2^8=3

OpenStudy (anonymous):

now u got it.........

OpenStudy (callisto):

Is that my fault again? I've interpreted your question twice and every time you said they are correct.. Now you're telling me that I'm wrong..

OpenStudy (callisto):

It is \[log_28 = log_22^3 = 3log_22 = 3\]

OpenStudy (callisto):

You've never told us what exactly the question is..

OpenStudy (callisto):

Can you scan us the question? I don't want to interpret it again.. I've been typing similar things for twice..

OpenStudy (anonymous):

ys this is the question

OpenStudy (anonymous):

i will draw

OpenStudy (anonymous):